# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ Appendix F More Related Works
Euclidean data clearly states which data points are associated with a particular individual, for example, tabular data. Unlike tabular data, graph association can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP, the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works (Gehrke et al., 2011; Mir and Wright, 2009; Wang et al., 2013; Chen et al., 2014; Lu and Miklau, 2014) started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider various node-privacy algorithms (Kasiviswanathan et al., 2013; Iftikhar and Wang, 2021; Day et al., 2016; Xia et al., 2021; Imola et al., 2021) tailored to specific graph statistics constrained by bounded node degrees. Node DP typically offers a stronger level of protection compared to edge privacy, as inclusion of a node implies the inclusion of some particular edges. Node DP is also considered harder to achieve than edge DP if maintaining reasonable utility loss.
Differentially private GNNs.
@@
Sharing model updates can lead to privacy risks, as adversaries may reconstruct training data. Such risks (Carlini et al., 2022; Ye et al., 2022) stem from overfitting, where neural networks memorize training data (Zhang et al., 2021). Graph structures, which encode relational information, are widely used in applications like intrusion detection (Apruzzese et al., 2023; Sommer and Paxson, 2010), social recommendation (Xia et al., 2023), and drug discovery (Zhong and Mottin, 2023). Graph neural networks (GNNs) have emerged as a key approach for learning over graph-structured data, but their message-passing mechanisms can leak sensitive information about nodes and their neighbors. To address these risks, several works (Sajadmanesh and Gatica-Perez, 2020, 2021; Sajadmanesh et al., 2023; Gupta et al., 2012) propose differentially private GNNs (Daigavane et al., 2021; Olatunji et al., 2021a; Sajadmanesh et al., 2023). This necessitates injecting calibrated noise after each message passing layer, where the noise scale is proportional to $K$ for a fixed privacy budget $\epsilon$. For instance, Wu et al. (Wu et al., 2022a) achieve edge-level privacy by adding Laplace noise to adjacency matrix entries under a small DP budget. Kolluri et al. (Kolluri et al., 2022) improve privacy-utility trade-offs by decoupling graph structure from the neural network architecture, querying the graph only when necessary. However, leveraging contractive hidden states in graph learning for enhanced privacy analysis remains an underexplored avenue.
Privacy analysis for iterations.
@@
Prior works on differentially private graph learning analyze cumulative privacy costs by composing iteration-wise privacy guarantees using calibrated noise and privacy composition theorems. These theorems enable modular analysis of complex algorithms by controlling the total privacy budget. Feldman et al. (Feldman et al., 2018, 2020) address privacy analysis for gradient computations over a single training epoch under smooth and convex loss functions. However, extending these guarantees to multiple epochs often relies on composition bounds, such as privacy amplification by subsampling. Recent advancements (Chourasia et al., 2021; Altschuler and Talwar, 2022; Altschuler et al., 2024) improve privacy analysis for convex and strongly convex losses by eliminating dependence on infinite iterations, leveraging the Rényi DP framework.
In the typical differentially private setting, each access to the graph data must be protected. For multi-hop message passing GNNs (e.g., GAP (Sajadmanesh et al., 2023)), the privacy guarantee scales linearly with the number of hops $K$. As $K$ increases, the enlarged noise scale and cumulative effect of noise can significantly degrade model utility—precisely when deeper architectures are needed to capture long-range dependencies. This creates a fundamental tension: models require sufficient depth to learn meaningful long-range interactions, yet each additional layer demands more privacy noise, eventually overwhelming the signal. Consequently, standard privacy mechanisms could fail to effectively learn in graph structures requiring multi-hop propagation, as the privacy-utility tradeoff becomes increasingly unfavorable with model depth. Despite progress, foundational questions about privacy costs remain challenging, even for smooth convex losses over bounded domains (Altschuler and Talwar, 2022). Recent works (Altschuler and Talwar, 2022; Ye and Shokri, 2022) reveal that privacy leakage does not increase indefinitely with iterations. They demonstrate that after a brief burn-in period, additional iterations of SGD do not significantly impact privacy.
Privacy Audit and Attacks.
The goal of privacy auditing (Steinke et al., 2023) is to empirically estimate the privacy leakage of an algorithm. The main techniques used for privacy auditing is typically mounting a membership inference attack (MIA) (Carlini et al., 2022; Olatunji et al., 2021b) that predicts the existence or absence of an individual example in the training set. In contrast to the upper bound obtained from theoretical privacy analysis, privacy auditing via MIA achieves the experimental true positive rate (TPR) and false positive rate (FPR) and then presents a lower bound on the privacy budget $\epsilon$.

### v2

@@ A-I1 Differential privacy for graph structures
Euclidean data clearly states which data points are associated with a particular individual, for example, tabular data. Unlike tabular data, graph association can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP, the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works [tcc/GehrkeLP11, icdm/MirW09, pakdd/WangWW13, vldb/ChenFYD14, kdd/LuM14] started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider various node-privacy algorithms [tcc/KasiviswanathanNRS13, pakdd/IftikharW21, sigmod/DayLL16, sigmod/XiaCKHT021, uss/ImolaMC21] tailored to specific graph statistics constrained by bounded node degrees. Node DP typically offers a stronger level of protection compared to edge privacy, as inclusion of a node implies the inclusion of some particular edges. Node DP is also considered harder to achieve than edge DP if maintaining reasonable utility loss.
A-I2 Differentially private GNNs
@@
Sharing model updates can lead to privacy risks, as adversaries may reconstruct training data. Such risks [sp/CarliniCN0TT22, ccs/0001MMBS22] stem from overfitting, where neural networks memorize training data [zhang2021understanding]. Graph structures, which encode relational information, are widely used in applications like intrusion detection [eurosp/ApruzzeseLS23, sp/SommerP10], social recommendation [icde/XiaSH0XP23], and drug discovery [kdd/ZhongM23]. Graph neural networks (GNNs) have emerged as a key approach for learning over graph-structured data, but their message-passing mechanisms can leak sensitive information about nodes and their neighbors. To address these risks, several works [arxiv/sajadmanesh2020differential, ccs/sajadmanesh2021locally, uss/sajadmanesh2023gap, tcc/GuptaRU12] propose differentially private GNNs [arxiv/DaigavaneNode21, corr/abs-2109-08907, uss/sajadmanesh2023gap]. This necessitates injecting calibrated noise after each message passing layer, where the noise scale is proportional to $K$ for a fixed privacy budget $\epsilon$. For instance, Wu et al. [sp/Wulinkteller22] achieve edge-level privacy by adding Laplace noise to adjacency matrix entries under a small DP budget. Kolluri et al. [ccs/KolluriBHS22] improve privacy-utility trade-offs by decoupling graph structure from the neural network architecture, querying the graph only when necessary. However, leveraging contractive hidden states in graph learning for enhanced privacy analysis remains an underexplored avenue.
A-I3 Privacy analysis for iterations
@@
Prior works analyze cumulative privacy costs by composing iteration-wise privacy guarantees using calibrated noise and privacy composition theorems. These theorems enable modular analysis of complex algorithms by controlling the total privacy budget. Feldman et al. [focs/FeldmanMTT18, stoc/FeldmanKT20] address privacy analysis for gradient computations over a single training epoch under smooth and convex loss functions. Recent advancements [nips/ChourasiaYS21, nips/AltschulerT22, siamcomp/AltschulerBT24] improve privacy analysis for convex and strongly convex losses by eliminating dependence on infinite iterations, leveraging the Rényi DP framework. While effective, Rényi DP is lossy. $f$-DP framework offers tighter analysis through hypothesis testing curves.
For multi-hop message passing GNNs (e.g., GAP [uss/sajadmanesh2023gap]), the privacy guarantee scales linearly with the number of hops $K$. Foundational questions about privacy costs remain challenging, even for smooth convex losses over bounded domains [nips/AltschulerT22]. Recent works [nips/AltschulerT22, nips/0001S22] reveal that privacy leakage does not increase indefinitely with iterations. They demonstrate that after a brief burn-in period, additional iterations of SGD do not significantly impact privacy.
A-I4 Overview of privacy attacks
Numerous privacy attacks have been developed against machine learning models, including membership inference, attribute inference, and reconstruction attacks. In the graph domain, these attacks are adapted to structured data, where the protected units are nodes, edges, or local subgraphs [uss/000100S022, sp/0011L0022, tpsisa/OlatunjiNK21]. MIA determine whether a specific node or edge appeared in the training graph, with variants studied in general ML models and specifically in GNNs. In $\mathsf{CARIBOU}$, privacy auditing module aims to empirically estimate the privacy leakage through an attacking algorithm [nips/0002NJ23, ccs/0001MMBS22]. Attribute inference attacks [olatunji2023does] aim to recover sensitive node or edge attributes from predictions. More recent graph-specific attacks further exploit message-passing behaviors to infer private structural information [uss/000100S022].
Attacking algorithms in $\mathsf{CARIBOU}$. The objective of $\mathsf{CARIBOU}$ is not to benchmark the full space of graph attacks, but to conduct a mechanism-level privacy audit of perturbed message passing under the black-box membership threat model [sp/CarliniCN0TT22, tpsisa/OlatunjiNK21] that underlies prior DP-GNN frameworks. In contrast to the upper bound obtained from theoretical privacy analysis, privacy auditing via MIA achieves the experimental true positive rate (TPR) and false positive rate (FPR) and then presents a lower bound on the privacy budget $\epsilon$ [sp/CarliniCN0TT22]. Within this setting, LinkTeller [sp/0011L0022] and the node-level GNN membership attack of Olatunji et al. [42] (G-MIA) are canonical choices: both are tailored to GNNs, operate in the transductive setting we consider, require only query access, and have open-source implementations. Using these two attacks provides a clean and reproducible evaluation of edge- and node-level leakage, enabling direct comparison with private GNNs.
Other attack families, such as attribute inference [olatunji2023does] or reconstruction attacks [zhang2022model, ijcai/ZhangLHWLLC21] require white-box knowledge. They aim to address different privacy notions or stronger threat models, and are therefore complementary rather than directly comparable to our membership-centric audit. Considering a practical and generic adversary, we do not include these attacks that require white-box access to internal embeddings or gradients, or that focus on DP notions such as attribute inference or graph-level reconstruction. Those attacks assume a strictly stronger adversary or target different sensitive objects than the edge- and node-level memberships that our DP guarantees protect. Extending our empirical evaluation to such attacks in other application scenarios is an interesting direction for future work.
Artifact Evaluation Appendix
A-A Artifact Summary
$\mathsf{CARIBOU}$ is a privacy-preserving GNN framework with a convergent privacy analysis, enabling deeper message passing under both edge-DP and node-DP without sacrificing utility. This artifact reproduces all experiments in Section VI, as summarized in Table XVI. It provides one-command scripts and pinned environments to regenerate: (i) privacy–utility results under EDP/NDP across nine datasets; (ii) $\epsilon$-curves, $K$-hops curves, $D$-curves, and heatmaps; and (iii) overhead measurements (time and memory). Overall, this artifact demonstrates that $\mathsf{CARIBOU}$ achieves superior privacy–utility balance and efficiency among private GNNs, with results verifiable through automated scripts and publicly available datasets.
TABLE XVI: Summary of Artifact & Evaluation Index.
Category Experiments Evaluation
PU PU-1 (Table III), PU-2 (Table XII), Sec. A-E1
PU-3 (Table XIII)
CRV CRV-$\epsilon$ (Figure 9), CRV-$K$ (Figure 10) Sec. A-E2
CRV-$D$, (Figure 8),CRV-H (Figure 11)
OV OV-E (Tables VII,IX), OV-N (Tables VIII,X) Sec. A-E3
“PU”: privacy-utility experiments; “CRV”: curves for accuracy as $\epsilon,K,D_{\max}$ varies; “AUD”: privacy auditing experiments; and “OV”: overhead measurements.
A-B Description & Requirements
This section provides all the information necessary to recreate the experimental setup to run our artifacts. All experiments can run on a commodity desktop machine.
A-B1 How to access
The’s main GitHub repository $\mathsf{CARIBOU}$ can be found at https://github.com/yuzhengcuhk/caribou-public, and the exact version of the code for the evaluation of artifacts is also available at https://doi.org/10.5281/zenodo.17539660 on the Zenodo platform. $\mathsf{CARIBOU}$ is released under the MIT License.
A-B2 Hardware dependencies
Our artifacts can be run on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. To ensure that all artifacts run correctly, it is recommended to use a machine with a similar configuration.
A-B3 Python environments
We provide a reproducible conda environment for evaluating all experiments, built on Python 3.9.20. Core system libraries are managed with conda, while most packages are pinned and installed via pip.
A-B4 Datasets
$\mathsf{CARIBOU}$ can be tested over nine datasets, including Photo and Computers [corr/abs-1811-05868], Cora and PubMed [sen2008collective], Facebook [corr/abs-1102-2166], Chain-S, Chain-M, Chain-L and Chain-X [nips/GuC0SG20].
A-B5 Artifact Architecture
core/ is the primary implementation directory, including:
  - methods/ # DP-GNN implementations
  - models/ # neural network architectures
  - modules/ #  GNN building blocks
  - privacy/ # privacy mechanisms
  - datasets/ # data loading
  - trainer/ # training and test loop
  - args/ # CLI parsing and configuration
  - utils.py # general utilities
A-C Artifact Installation & Configuration
A-C1 Download the code
To download the artifact, we recommend cloning via HTTPS. Execute the following commands to fetch the repository, and then enter the $\mathsf{CARIBOU}$’s directory.
1git clone https://github.com/yuzhengcuhk/caribouCaemra-public.git
2cd caribou-public/
A-C2 Build the environment
Inside caribou-public/, we provide a script setup_minimal_env.sh to quickly and conveniently build $\mathsf{CARIBOU}$’s environment for evaluators. The shell script setup_minimal_env.sh creates a conda environment (caribou-minimal) installs the core libraries from the CUDA 11.7 wheel index.
1chmod +x ./setup_minimal_env.sh
2./setup_minimal_env.sh
3conda activate caribou-minimal
4python train.py mlp-dp --dataset cora --epsilon 2
For users who prefer pip-based installs, we also provide requirements_minimal_caribou.txt that lists the same minimal set and helps evaluators to install relevant packages. After the environment is successfully provisioned, execute the example run via the third command. This launches a DP-MLP training sanity check and prints the results.
A-C3 train.py entry point
train.py serves as the unified launcher for all experiments. The --method flag selects the pipeline family and its privacy regime, including $\mathsf{CARIBOU}$ and all baseline works in this paper: mlp, mlp-dp, gap-inf, gap-edp, gap-ndp, sage-inf, sage-edp, sage-ndp, dpdgc-inf, dpdgc-edp, dpdgc-ndp, caribou-inf, caribou-edp, caribou-ndp. Internally, the entry point dispatches to the corresponding implementation in core/methods/.
A-D Major Claims
  • (C1): $\mathsf{CARIBOU}$ can outperform all other baselines in most cases with varying $\epsilon$, no matter whether it is EDP or NDP. $\mathsf{CARIBOU}$ achieves a more favorable privacy-utility trade-off.
  • (C2): $\mathsf{CARIBOU}$ consumes relatively low computational overhead in execution time and max memory usage.
A-E Evaluation
This section enumerates the end-to-end steps and experiments required to evaluate our artifact and validate the paper’s claims (C1, C2 in Section A-D) for $\mathsf{CARIBOU}$. The full evaluation can be completed in approximately 20 human minutes and about 8 compute hours, subject to hardware. We assume the host machine is correctly configured with the dependencies described in Section A-C. The same instructions can also be found in the caribou-public/, i.e., top-level README to provide a consistent and reproducible workflow.
A-E1 Evaluation (E1)
[PU-1 (5 human-minutes, 3 computer-hours) + PU-2 (5 human-minutes, 3 computer-hours) + PU-3 (5 human-minutes, 3 computer-hours)]
Under AE/PU/, we provide two scripts for quickly reproducing privacy–utility experiments. run_scripts_all.sh executes the full experiments across the configured datasets and privacy budgets to regenerate all three accuracy tables (Table III, Table XII, Table XIII) reported in the paper. It is intended for a complete reproduction run on a machine with sufficient compute. For convenience to evaluators, run_scripts_computers.sh is a targeted script that launches the exactly same pipelines on the computers dataset, serving as a quick sanity check to verify accuracy results. To reduce the workload, we only compare $\mathsf{CARIBOU}$ with the generally strongest baseline GAP in artifact evaluation. It creates an output directory (AE_outputs/PU/) for storing results. Within the caribou_public directory, execute:
1chmod +x AE/PU/pu1_run_scripts_computers.sh
2./AE/PU/pu1_run_scripts_computers.sh # PU-1
3
4chmod +x AE/PU/pu2_run_scripts_computers.sh
5./AE/PU/pu2_run_scripts_computers.sh # PU-2
6
7chmod +x AE/PU/pu3_run_scripts_computers.sh
8./AE/PU/pu3_run_scripts_computers.sh # PU-3
Each evaluation item (i.e., each table entry) emits a plain-text report. Evaluators can open the corresponding .txt file to read either the top accuracy or the mean accuracy over three independent runs. Because training seeds are randomized and i.i.d noise is sampled every time, some fluctuations across executions are expected. The Chain datasets are particularly noise–sensitive, and thus can exhibit comparatively larger variation. The Chain datasets are particularly noise-sensitive and thus can exhibit comparatively larger variation. This sensitivity is due to their structure: non-zero features are present only at the first node of each chain. Information must propagate from this source, and it can be degraded by noise accumulation during propagation. This task is already challenging with small number of nodes, and its difficulty can be unwittingly amplified by the random train/val/test split. To keep runtime modest as required by artifact evaluation, the provided scripts sweep only a subset of the hyper-parameters used for the paper’s final tables. As a result, the top accuracy observed by evaluators during reproduction sometimes may be lower than the maxima reported in the paper. This is expected as the paper results reflect a broader search grid. Despite these effects, $\mathsf{CARIBOU}$ exceeds the accuracy of competing baselines in the majority settings (Claim C1).
A-E2 Evaluation (E2)
[CRV-$\epsilon$ (5 human-minutes, 0.5 computer-hours) + CRV-$K$ (5 human-minutes, 0.5 computer-hours) + CRV-$D$ (5 human-minutes, 0.5 computer-hours) + + CRV-H (5 human-minutes, 0.5 computer-hours)] Before re-generating Figures 9,10,8,11 we need to install jupyter and register this environment as a Jupyter kernel. Notably, each of the following commands is a single line. Some PDF viewers soft-wrap lines, which can corrupt copy-and-paste. To prevent errors, paste each command into a plain-text (.txt) file, remove any spurious whitespace, then paste and execute it in the terminal. To reproduce all figures, execute the following commands within the caribou_public directory:
1python -m pip install jupyter
2python -m pip install seaborn==0.13.2
3
4python -m ipykernel install --user --name caribou-minimal --display-name "Python(caribou-minimal)"
5
6jupyter nbconvert --to notebook --execute --inplace --ExecutePreprocessor.kernel_name=caribou-minimal ./AE/CRV/eps_hop_plots.ipynb # CRV-eps, CRV-K
7
8jupyter nbconvert --to notebook --execute --inplace --ExecutePreprocessor.kernel_name=caribou-minimal ./AE/CRV/degree_plots.ipynb # CRV-D
9
10jupyter nbconvert --to notebook --execute --inplace --ExecutePreprocessor.kernel_name=caribou-minimal ./AE/CRV/heatmap.ipynb # CRV-H
The evaluated results are embedded within the corresponding .ipynb notebooks and can be inspected directly after execution. In addition, the generated figures are stored in AE_outputs/CRV/.
A-E3 Evaluation (E3)
[OV-E (5 human-minutes, 2 computer-hours) + OV-N (5 human-minutes, 2 computer-hours)] We provided two scripts to execute the overhead benchmarks for EDP and NDP, respectively. Within the caribou_public directory, execute:
1chmod +x AE/OV/ove_run_scripts.sh
2./AE/OV/ove_run_scripts.sh
3
4chmod +x AE/OV/ovn_run_scripts.sh
5./AE/OV/ovn_run_scripts.sh
Upon completion of each run, the corresponding computational overhead results are available in AE_outputs/OV/. This directory contains the actual execution time and the maximum memory usage results referenced in Tables VII,IX,VIII,X. Note that absolute values are hardware–dependent; therefore, the numbers obtained by the evaluators may differ from those exactly reported in the paper. This variation is expected. Provided the machine offers sufficient compute resources, the qualitative trends and relative comparisons should remain consistent with our results (Claim C2).